OpenStudy (anonymous):

If you want to represent the 4.5x10^9 years of geologic time with a roll of tape so that 1.00 inch represents 1.00x10^6 years. how long would the tape have to be in miles?

OpenStudy (kropot72):

The length in inches is\[\frac{4.5\times 10^{9}}{1.00\times 10^{6}}\] What do you get for this?